OpenStudy (anonymous):

2. A prism has total surface area of 360 m^2 and volume of 60 m^3. If the length, width, and height are reduced to half their original sizes, what will be the following? a.) the new surface area b.) the new volume If the length, width, and height of the original are increased to triple their original sizes, what will be the following? c.) the new surface area d.) the new volume Show your calculations for all four parts.

Directrix (directrix):

A prism has total surface area of 360 m^2. the length, width, and height are reduced to half their original sizes, what will be the new surface area? ------------------ Theorem Needed: If two solids are similar, the square of the scale factor of the two solids is equal to the ratio of any two corresponding area measurements of the solids.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

1440?

Directrix (directrix):

1440? I don't think so. The formula is upside down.

OpenStudy (anonymous):

60 over V = 1 over 8?

Directrix (directrix):

The values should be getting smaller. Back to surface area. sing the theorem, the old surface area is to the new surface area as the square of the scale factor. 360/x = (2/1)^2 The original has dimensions twice the second one. 360/x = 4/1 Cross multiply and solve for x. Now, what is this?

OpenStudy (anonymous):

90?

Directrix (directrix):

That is what I got.

Directrix (directrix):

Now to volume.

Directrix (directrix):

60/V = (2/1)^3 60/V = 8/1 Cross multiply and solve for V.

Directrix (directrix):

Answer?

OpenStudy (anonymous):

7.5

Directrix (directrix):

60/8 = 7.5 cubic meters

OpenStudy (anonymous):

so my answer was wrong?

Directrix (directrix):

I thought we got the same thing. 7.5

Directrix (directrix):

Part II: Here are the two theorems we are using: If two solids are similar, the cube of the scale factor of the two solids is equal to the ratio of the volumes.   If two solids are similar, the square of the scale factor of the two solids is equal to the ratio of any two corresponding area measurements of the solids.

OpenStudy (anonymous):

no i was just making sure xD

OpenStudy (anonymous):

sorry xd moving on

Directrix (directrix):

Okay. This time the original solid has its dimensions tripled. The scale factor from small to large is 1 to 3. 360/S = (1/3)^2 where S is the surface area of the larger solid

OpenStudy (anonymous):

9720?

Directrix (directrix):

I messed up again. 1/3 times 1/3 = 1/9, not 1/27.

Directrix (directrix):

360/S = 1/9 Cross multiply and solve for S.

OpenStudy (anonymous):

3240?

Directrix (directrix):

3240 square meters is what I got. ------- 60/V = (1/3)^3 where V is the volume of the larger solid. 60/V = 1/27 Cross multiply and solve for V.

OpenStudy (anonymous):

1620

Directrix (directrix):

That's what I got.

OpenStudy (anonymous):

so that's all ?

OpenStudy (anonymous):

k last question

Directrix (directrix):

Put it in a new thread.

Directrix (directrix):

Oh, that was the last question. I thought you were saying that there is another.

OpenStudy (anonymous):

there is i just asked it